OpenStudy (anonymous):

Find the sum to n terms \[\frac{ 1 }{ 1*2 }+\frac{ 1 }{ 2*3 }+\frac{ 1 }{ 3*4 }\]

OpenStudy (experimentx):

find the general terms ... make partial fractions .. then use telescoping.

OpenStudy (anonymous):

WAT IS TELESCOPING

OpenStudy (experimentx):

first do the first two steps

OpenStudy (anonymous):

i got \[\frac{ 1 }{ n(n+1) }\]

OpenStudy (experimentx):

then do the second step.

OpenStudy (anonymous):

Hw to do that

OpenStudy (experimentx):

don't you know about partial fractions?

OpenStudy (anonymous):

\[\frac{ A }{ n }+\frac{ b }{ n+1 }\] like this

OpenStudy (experimentx):

yeah .. A=1 and b=-1

OpenStudy (anonymous):

\[\frac{ 1 }{ n }-\frac{ 1 }{ n+1 }\]

OpenStudy (experimentx):

better google for video on Heaviside cover up method for partial fractions. it's not easier.

OpenStudy (experimentx):

not -> *lot

OpenStudy (anonymous):

i knw....partial fraction

OpenStudy (anonymous):

then nw

OpenStudy (experimentx):

add the series.

OpenStudy (anonymous):

Sn = n(n+1)/2

OpenStudy (experimentx):

that's harmonic series ... you can't do that.

OpenStudy (anonymous):

Oh..... Sn = a1+a2........an

OpenStudy (experimentx):

put n=1, 2, 3, ... and evaluate the sum.

OpenStudy (anonymous):

a1 = 1/1 - 1/2 a2=1/2-1/3.......................... an = 1/n -1/n+1

OpenStudy (experimentx):

yeah yea ... put those up and cancel the + - and get your final result 1 - 1/n+1

OpenStudy (anonymous):

1/1 - 1/n+1 will remain

OpenStudy (anonymous):

Yup..) thxxx..)

OpenStudy (anonymous):

\[\frac{ 1 }{ 1\times2 }+\frac{ 1 }{ 2\times3 }+\frac{ 1 }{ 3\times4 }+......\]\[=\frac{ 2-1 }{ 1\times2 }+\frac{ 3-2}{ 2\times3 }+\frac{ 4-3 }{ 3\times4 }+......\]\[=1-\frac{ 1 }{ 2 }+\frac{ 1 }{ 2 }-\frac{ 1 }{ 3 }+\frac{ 1 }{ 3}-\frac{ 1 }{ 4 }+\frac{ 1 }{ 4}-\]\[=1\]

OpenStudy (anonymous):

It is to infty.